What do the following mean if:
a.) an expression is a function
b.) y is directly proportional to x (give an example)
c.) the coefficient of the third term is 5 ( give an example)
d.) a function only has one root
e.) a function is a polynomial
can someone help me with explanations for the above statements please? THANKS!
any help will be nice.

Answer by stanbon(29603)

a.) an expression is a function
y=3x+2 is an equation that is a function because for each
value of x there is exactly one value of y. 3x+2 is an
expression; it is not an equation and not a function.
----------------------------------------------
b.) y is directly proportional to x (give an example)
y=kx means y and x are directly proportional;
Example: If 4=k(8) then 10=k(20) In both cases k=1/2
---------------------------------------------------
c.) the coefficient of the third term is 5 ( give an example)
If y=2x^3+3x^2+4x-8 the 3rd term is 4x
-------------------------------------------
d.) a function only has one root
If y=3x+2 and we let y=0 the value of x will be -2/3 and
that value of x is called a root or zero of the function.
-----------------------------------------
e.) a function is a polynomial
A polynomial is an algebraic expression comprising more
than one term. A function is a special equation; it is
never called a polynomial.
-----------------------------
Cheers,
Stan H.